OpenStudy (anonymous):

simplify the expression 3x9-4+6\2squared

OpenStudy (anonymous):

\[3\times9-4+6\div 2^{2}\]

OpenStudy (anonymous):

first you want to start here 2^2 = 4 that would give you 3 X 9 - 4 + 4

OpenStudy (anonymous):

\[3 \times 9 - 4 + 6 \div 4\]

OpenStudy (anonymous):

\[ 3\times9-4+6\div4\] yeah i had that part down and when i divide 6/4 i get 1.5 then what?

OpenStudy (anonymous):

mulitply 3 x 9

OpenStudy (anonymous):

i already did that now its 27-4+1.5 then i know you subtract 4 from 27 its 23+1.5 but out of the answers to choose from they are all fractions what i dont get is how it turns into a fraction

OpenStudy (anonymous):

so 24.5 = 49/2

OpenStudy (anonymous):

oh what you could do is use 1/2 instead of .5

OpenStudy (anonymous):

how did you get 49/2

OpenStudy (anonymous):



OpenStudy (anonymous):

ok i got it now

OpenStudy (anonymous):

i just used my calculator to get the fraction